Data from a representative sample were used to estimate that 32% of all computer users in a recent year had tried to get on a Wi-Fi network that was not their own in order to save money. You decide to conduct a survey to estimate this proportion for the current year. What is the required sample size if you want to estimate this proportion with a margin of error of 0.04? Calculate the required sample size first using 0.32 as a preliminary estimate of p.


1
Expert's answer
2021-05-31T01:41:56-0400

Margin of error:

"ME=Z_{\alpha}\sqrt{\frac{p(1-p)}{n}}=0.04"

where n is sample size,

p is estimated proportion.

Then:

"n=\frac{Z^2_{\alpha }p(1-p)}{ME^2}"


For example, for confidence interval 95%:

"Z_{0.95}=1.96"


"n=\frac{1.96^2\cdot0.32(1-0.32)}{0.04^2}=522.45"


So, the required sample size:

"n=523"